# TheoremExplainAgent: Towards Multimodal Explanations for LLM Theorem Understanding

### v1

@@ Abstract
TheoremExplainAgent: Towards Multimodal Explanations for LLM Theorem Understanding
@@ 1 Introduction
Our experiments with TheoremExplainAgent yielded both promising results and clear areas for improvement in AI-generated multimodal theorem explanations. On the positive side, a key achievement was the system’s ability to generate extended video explanations, reaching durations of up to 10 minutes. This represents a significant advancement over agentless approaches, which we found to be limited to approximately 20-second videos. Furthermore, TheoremExplainAgent demonstrated versatility across different STEM disciplines, successfully creating videos for Mathematics, Physics, Chemistry, and Computer Science. Importantly, we observed that video-based theorem explanations inherently expose deeper reasoning flaws in AI systems that text-based evaluations often miss. Unlike text-based multiple-choice questions, where models can exploit superficial cues, generating visual-theorem explanations necessitates that the AI explicitly encodes structural and procedural knowledge, thus making underlying errors more apparent. In particular, the o3-mini model exhibited robust performance at varying levels of theorem difficulty, indicating a capacity to handle both fundamental and complex concepts. However, despite these successes, limitations persist. While the system could generate textually accurate explanations, the visual quality and pedagogical structure of the videos require further refinement. Generated animations frequently exhibited minor visual layout inaccuracies, such as misaligned text elements, overlapping shapes, and inconsistent object sizes. These visual errors, though often subtle, became more pronounced and potentially distracting, particularly in the medium and hard difficulty levels of our TheoremExplainBench.
@@ 2.1 LLM and Agents
The rapid advancements in large language models (LLMs) and large vision-language models (VLMs) have unlocked unprecedented capabilities in understanding multimodal content. Models such as GPT-4 OpenAI (2023), Gemini Gemini-Team et al. (2024), Claude-3.5 Sonnet v1 Anthropic (2024), and DeepSeek DeepSeek-AI et al. (2024) have demonstrated strong abilities in processing complex textual information and analyzing visual inputs within a unified framework Zhang et al. (2023b). These breakthroughs have enabled transformative applications across various domains, including visual content understanding Hu et al. (2023); Ku et al. (2023), code generation Nijkamp et al. (2023); Jimenez et al. (2024); Yang et al. (2024a), and reasoning over structured data. To tackle complex tasks, researchers have explored LLM agents: AI systems that leverage LLMs to autonomously reason, plan, and execute tasks by interacting with structured environments or external tools. These agents have been deployed in various goal-oriented applications, such as scientific discovery Lu et al. (2024); Si et al. (2024); Schmidgall et al. (2025), coding solutions Abramovich et al. (2024), multimodal visual generation He et al. (2024), and computer environment interaction Xie et al. (2024). In this work, we extend the use of LLM agents into the domain of theorem explanation and visualization.
@@ 3.2 TheoremExplainAgent (TEA)
We develop TheoremExplainAgent (TEA), an agentic pipeline designed to automate the generation of videos using multiple specialized agents as shown in Figure 3. The process begins with the planner agent, which creates a high-level video plan according to the specified theorem. This plan consists of multiple scenes, each corresponding to a key segment of the resulting video. Once the initial plan is created, the planner agent refines the details of each scene, breaking them down into smaller components that define the specific visual elements, animations, and transitions needed. These detailed scene descriptions are then passed to the coding agent, which generates the corresponding Python code. The voiceover is also generated through a text-to-speech service. Finally, the Python scripts are executed to produce the final video, which reflects the narrative or instructional goals outlined in the video plan. If the generated Python code encounters an error, the coding agent will review the error and generate a revised version of the code. We set a maximum of $N$ attempts where $N=5$. If this limit is exceeded, we mark the generation as unsuccessful.
@@ 4 Experimental Results
For the agent candidates in TheoremExplainAgent, we experimented with GPT-4o OpenAI (2023), Gemini 2.0 Flash DeepMind (2025), Claude 3.5 v1 Anthropic (2024), and o3-mini OpenAI (2025). Each candidate was used for both the planner agent and coding agent, ensuring consistency across configurations. We evaluated all agents across 240 theorems from TheoremExplainBench, comparing their performance under different setups. Our findings indicate that an agentless approach fails to generate videos longer than 20 seconds, whereas TheoremExplainAgent successfully produces videos of up to 10 minutes. Consequently, all experimental results presented below are based on the agentic approach.
@@
Our experiments with the RAG setup yielded mixed results, as shown in Table 1 and Table 2. While RAG was expected to enhance function understanding and streamline object construction, its effectiveness proved inconsistent. Although retrieving documentation and code examples provided additional context, the results often misaligned with specific use cases. Many retrieved references were too generic or lacked relevance, leading to incorrect function calls and suboptimal parameter choices. These findings align with previous research emphasizing that retrieval quality is crucial. Poorly structured documentation and imprecise retrieval can significantly reduce the effectiveness of RAG-based approaches Soman and Roychowdhury (2024).
@@ 4.3 Case Study
We included representative video outputs in Figure 6. This figure demonstrates that TheoremExplainAgent is capable of generating high-quality exploratory videos. For example, in Mathematics, the model effectively visualizes concepts such as Riemann sums, using animated grids and function plots to illustrate integral approximations. In Chemistry, the system successfully explains the Octet Rule, leveraging atomic models to depict electron sharing and bonding interactions. In Physics, it generates electromagnetic wave simulations, showcasing wave propagation and spectral analysis. In Computer Science, it produces a clear demonstration of Run-Length Encoding, using side-by-side comparisons of raw and compressed data representations. We examined more generated videos carefully and observed that videos in Mathematics, Physics, and Computer Science generally exhibit higher visual quality and coherence compared to those in Chemistry. One notable observation is that Chemistry-related visualizations often rely on simple geometric primitives to depict complex lab apparatus and molecular structures, which can limit their clarity and effectiveness. Additionally, most of the generated videos exhibit minor element layout issues, such as overlapping texts, inconsistent sizes, or suboptimal object positioning, which slightly affects the overall presentation quality, as illustrated in Figure 7.
We also found that visual explanations more effectively reveal reasoning errors than text, facilitating error diagnosis. From Figure 5, we observe that while the text-based explanation allows us to detect that the model’s answer is incorrect, it does not provide insight into why the mistake occurred. It seems the model understand the chain code theorem, but it applies it incorrectly. Such explanation is making it difficult to pinpoint the exact reasoning flaw. In contrast, the video-based explanation clearly exposes the model’s misunderstanding, as incorrect movement direction encodes and misplaced arrows reveal how the model misinterpreted the chain coding process. This demonstrates that visual explanations not only confirm incorrect reasoning but also uncover the underlying cause of errors, making them a more effective diagnostic tool for analyzing AI-generated outputs.
@@ 5 Conclusion
This paper introduces TheoremExplainAgent, a novel agentic approach for generating multimodal theorem explanations through structured video content. Our study demonstrates that integrating visual explanations significantly enhances the clarity and interpretability of theorem reasoning, surpassing text-based methods alone. To systematically evaluate AI-generated explanations, we present a benchmark spanning multiple disciplines with five automated evaluation metrics. Our experiments reveal that agentic planning is crucial for producing long-form, coherent explanations, with o3-mini achieving the highest success rate and overall performance. However, challenges remain in visual element layout, emphasizing the need for improved spatial reasoning and refinement in AI-generated animations. Additionally, our findings underscore the importance of multimodal explanations in identifying reasoning flaws that text-based assessments often miss, reinforcing the role of structured visual communication in AI-driven theorem understanding. Looking ahead, future work should focus on enhancing visual structuring techniques, improving agent coordination, and advancing video understanding to further refine multimodal explanations for LLM-driven theorem comprehension.
@@ 6 Limitations
While our approach demonstrates the potential of AI-generated multimodal theorem explanations, several limitations remain. AI models still struggle with complex visual structuring, particularly in consistent elements layout in long-form explanations. Retrieval-augmented generation (RAG) also requires more tokens, increasing computational costs and inference time, which may impact scalability.
@@ Appendix A Gallery
Figure 6: We show the high-quality videos generated by TheoremExplainAgent,across the four STEM domains.
Figure 7: We show the poorly generated videos from TheoremExplainAgent, zooming in the artifacts.
Figure 8: Comparison on a scene of a high quality animation and a low quality animation.
Appendix B Prompt Templates
@@
Appendix C Supplementary Information
C.1 Human Annotation Process
We recruited 12 student volunteers in our annotation process. We explained to the annotators that their annotations were to be used in our study only and would not be released publicly.
Figure 9: The user interface of our annoatation website.
C.2 Runtime Statistics
Agent Input Tokens Output Tokens Cost(USD) Time(s)
GPT-4o 350000 84000 1.71 1120
GPT-4o + RAG 840000 84000 2.94 1260
Claude 3.5-Sonnet v1 350000 91000 2.42 2240
Claude 3.5-Sonnet v1 + RAG 1050000 101500 4.67 2380
Gemini 2.0-Flash 595000 119000 0.1 1120
Gemini 2.0-Flash + RAG 1120000 119000 0.16 1260
o3-mini (medium) 434000 154000 1.16 1680
o3-mini (medium) + RAG 945000 154000 1.72 1820
C.3 Potentials for Future Research

### v2

# TheoremExplainAgent: Towards Video-based Multimodal Explanations for LLM Theorem Understanding

@@ Abstract
TheoremExplainAgent: Towards Video-based Multimodal Explanations for LLM Theorem Understanding
@@ 1 Introduction
Our experiments with TheoremExplainAgent yielded both promising results and clear areas for improvement in AI-generated multimodal theorem explanations. On the positive side, a key achievement was the system’s ability to generate extended video explanations, reaching durations of up to 10 minutes. This represents a significant advancement over agentless approaches, which we found to be limited to approximately 20-second videos. Furthermore, TheoremExplainAgent demonstrated versatility across different STEM disciplines, successfully creating videos for Mathematics, Physics, Chemistry, and Computer Science. Importantly, we observed that video-based theorem explanations inherently expose deeper reasoning flaws in AI systems that text-based evaluations often miss. Unlike text-based multiple-choice questions, where models can exploit superficial cues, generating visual-theorem explanations necessitates that the AI explicitly encodes structural and procedural knowledge, thus making underlying errors more apparent. In particular, the o3-mini model exhibited robust performance at varying levels of theorem difficulty, indicating a capacity to handle both fundamental and complex concepts. However, despite these successes, limitations persist. While the system could generate textually accurate explanations, the visual quality and pedagogical structure of the videos require further refinement. Generated animations frequently exhibited minor visual layout inaccuracies, such as misaligned text elements, overlapping shapes, and inconsistent object sizes. These visual errors, though often subtle, became more pronounced and potentially distracting. This happens particularly in the medium and hard difficulty levels of our TheoremExplainBench.
@@ 2.1 LLM and Agents
The rapid advancements in large language models (LLMs) and large vision-language models (VLMs) have unlocked unprecedented capabilities in understanding multimodal content. Models such as GPT-4 OpenAI (2023), Gemini Gemini-Team et al. (2024), Claude-3.5 Sonnet v1 Anthropic (2024), and DeepSeek DeepSeek-AI et al. (2024) have demonstrated strong abilities in processing complex textual information and analyzing visual inputs within a unified framework Zhang et al. (2023b). These breakthroughs have enabled transformative applications across various domains, including visual content understanding Hu et al. (2023); Ku et al. (2023), code generation Nijkamp et al. (2023); Jimenez et al. (2024); Yang et al. (2024a), and reasoning over structured data. To tackle complex tasks, researchers have explored LLM agents: AI systems that leverage LLMs to autonomously reason, plan, and execute tasks by interacting with structured environments or external tools. These agents have been deployed in various goal-oriented applications, such as scientific discovery Lu et al. (2024); Si et al. (2024); Schmidgall et al. (2025), coding solutions Abramovich et al. (2024), multimodal visual generation He et al. (2024a), and computer environment interaction Xie et al. (2024). In this work, we extend the use of LLM agents into the domain of theorem explanation and visualization.
@@ 3.2 TheoremExplainAgent (TEA)
We develop TheoremExplainAgent (TEA), an agentic pipeline designed to automate the generation of videos using multiple specialized agents as shown in Figure 3. The process begins with the planner agent, which creates a high-level video plan according to the specified theorem. This plan consists of multiple scenes, each corresponding to a key segment of the resulting video. Once the initial plan is created, the planner agent refines the details of each scene, breaking them down into smaller components that define the specific visual elements, animations, and transitions needed. These detailed scene descriptions are then passed to the coding agent, which generates the corresponding Python code. The voiceover is also generated through a text-to-speech service. Finally, the Python scripts are executed to produce the final video, which reflects the narrative or instructional goals outlined in the video plan. If the generated Python code encounters an error, the coding agent will review the error and generate a revised version of the code. We set a maximum of $N$ attempts where $N=5$. If this limit is exceeded, we mark the generation as unsuccessful. We found that at $N=0$, the success rate is extremely low while $N=5$ achieves up to 90% success rate, as discussed in Table 3.
@@
Category $N=0$ $N=1$ $N=2$ $N=3$ $N=4$ $N=5$
Difficulty: Easy 7% / 5% 51% / 33% 73% / 66% 86% / 71% 91% / 73% 93% / 73%
Difficulty: Medium 0% / 5% 33% / 43% 75% / 66% 83% / 72% 88% / 76% 91% / 77%
Difficulty: Hard 3% / 3% 46% / 35% 81% / 51% 90% / 68% 95% / 71% 96% / 73%
Table 3: Combined cumulative theorem success rates (Baseline / RAG) for o3-mini (medium) with varying N attempts. The overall success rate significantly improves with more attempts, reaching its peak at $N=5$.
@@ 4 Experimental Results
For the agent candidates in TheoremExplainAgent, we experimented with GPT-4o OpenAI (2023), Gemini 2.0 Flash DeepMind (2025), Claude 3.5 v1 Anthropic (2024), and o3-mini OpenAI (2025). Each candidate was used for both the planner agent and coding agent, ensuring consistency across configurations. We evaluated all agents across 240 theorems from TheoremExplainBench, comparing their performance under different setups. Our findings indicate that an agentless approach fails to generate videos longer than 20 seconds, whereas TheoremExplainAgent successfully produces videos of up to 10 minutes. Consequently, all experimental results presented in this paper are based on the agentic approach.
@@
Our experiments with the RAG setup yielded mixed results, as shown in Table 1 and Table 2. While RAG was intended to enhance function understanding and streamline object construction, its effectiveness proved inconsistent. Although retrieval of documentation and code examples provided additional context, the results often misaligned with specific use cases. Many retrieved references were overly generic or lacked relevance, leading to incorrect function calls and suboptimal parameter choices. These findings are consistent with previous research highlighting the critical importance of retrieval quality. Poorly structured documentation and imprecise retrieval can significantly compromise the effectiveness of RAG-based approaches Soman and Roychowdhury (2024).
[…]
@@ 4.1 Correlation Study
4.2 Interpretability Study
We found that visual explanations more effectively reveal reasoning errors than text, facilitating error diagnosis. From Figure 5, we observe that while the text-based explanation allows us to detect that the model’s answer is incorrect, it does not provide insight into why the mistake occurred. It seems the model understand the chain code theorem, but it applies it incorrectly. Such explanation is making it difficult to pinpoint the exact reasoning flaw. In contrast, the video-based explanation clearly exposes the model’s misunderstanding, as incorrect movement direction encodes and misplaced arrows reveal how the model misinterpreted the chain coding process. This demonstrates that visual explanations not only confirm incorrect reasoning but also uncover the underlying cause of errors, making them a more effective diagnostic tool for analyzing AI-generated outputs.
To quantitatively assess whether video explanations better highlight LLM reasoning problems, we designed a human study. 15 participants were first shown a textual explanation of a theorem containing a subtle reasoning flaw and asked to judge its correctness. They were then shown a video explanation of the same theorem, embodying the same reasoning flaw, and asked to re-evaluate. Participants also rated the intuitiveness of both explanations on a scale of 1 to 5. Initially, all 15 participants judged the textual explanation as correct. After watching the video, 9 participants (60%) revised their judgment to "incorrect," successfully identifying the conceptual flaw that was only apparent through the visual narration. The average intuitiveness rating improved from 3.3 for the textual explanation to 3.9 for the video explanation. This suggests that multimodal explanations can be more effective in revealing and helping users understand reasoning errors.
@@ 4.3 Error Analysis
Figure 6: Video quality comparison by video frame between LTXVideo, Veo2 and TheoremExplainAgent, where $t$ represents the proportion of the total video duration. $t$ correspond to the time. The approximate number of $T$ for LTXVideo, Veo2 and TheoremExplainAgent are 7 seconds, 8 seconds, and 5 minutes respectively.
@@ 4.4 Case Study
We included representative video outputs in Figure 7, demonstrating that TheoremExplainAgent is capable of generating high-quality exploratory videos. For example, in Mathematics, the model effectively visualizes concepts such as Riemann sums, using animated grids and function plots to illustrate integral approximations. In Chemistry, the system successfully explains the Octet Rule, leveraging atomic models to depict electron sharing and bonding interactions. In Physics, it generates electromagnetic wave simulations, showcasing wave propagation and spectral analysis. In Computer Science, it produces a clear demonstration of Run-Length Encoding, using side-by-side comparisons of raw and compressed data representations. Videos in Mathematics, Physics, and Computer Science typically show higher visual quality and coherence than those in Chemistry. One notable observation is that Chemistry-related visualizations often rely on simple geometric primitives to illustrate complex lab equipment and molecules, which can limit their clarity and effectiveness. Additionally, most of the generated videos exhibit minor element layout issues, such as overlapping texts, inconsistent sizes, or suboptimal object positioning, which slightly affects the overall presentation quality, as illustrated in Figure 8.
4.5 Text-to-Video Model Baselines
To access whether explicit reasoning capabilities provided by LLM-based agents are essential for generating coherent theorem explanations, we conducted a baseline comparison using non-LLM-based text-to-video models. Specifically, we examined the performance of the recent open-source model LTXVideo HaCohen et al. (2024) and the closed-source model Veo2 Veo-Team et al. (2024). For this comparison, we prompted each model with: “a Manim-style explanatory video explaining <theorem>” for 20 randomly selected theorems from TheoremExplainBench. As illustrated in Figure 6, the resulting videos were frequently visually incoherent, often manifesting as random noise or lacking meaningful relation to the intended scientific content. These outputs also lacked the structured, pedagogical qualities necessary for effective theorem explanations. Moreover, these models do not include voiceover narration, an integral component of our multimodal output. This highlights that while text-to-video models can synthesize visual content from text, they lack the explicit reasoning capabilities required to generate domain-specific explanatory videos. In contrast, TheoremExplainAgent enables the creation of detailed, pedagogically sound videos that align with the theorems.
@@ 5 Conclusion
This paper introduces TheoremExplainAgent, a novel agentic approach for generating multimodal theorem explanations through structured video content. We demonstrates that integrating visual explanations significantly enhances the clarity and interpretability of theorem reasoning, surpassing text-based methods alone. We also present a benchmark spanning multiple disciplines with five automated evaluation metrics. Experiments reveal that agentic planning is crucial for producing long-form, coherent explanations, with o3-mini achieving the highest success rate and overall performance. However, challenges remain in visual element layout, emphasizing the need for improved spatial reasoning and refinement in AI-generated animations. Additionally, our findings underscore the importance of multimodal explanations in identifying reasoning flaws that text-based assessments often miss, reinforcing the role of structured visual communication in AI-driven theorem understanding.
@@ 6 Limitations
While our approach demonstrates the potential of AI-generated multimodal theorem explanations, several limitations remain.
Visual Structuring Challenges.
TheoremExplainAgent exhibits limitations in visual structuring, with issues such as misaligned text, overlapping shapes, and inconsistent sizes. These visual imperfections, though sometimes subtle, can distract from the intended educational value and hinder comprehension, particularly in complex topics like Chemistry. As mentioned in Section 4.4 and illustrated in Table 2, both human and automated evaluations identified shortcomings in element layout and visual consistency, highlighting the need for further refinement in visual design.
Inconsistent Retrieval-Augmented Generation.
The reliance on retrieval-augmented generation (RAG) to support Manim code generation proved inconsistent in practice. As noted in Section 4.4 and shown in Table 5, retrieved code snippets were often irrelevant or overly generic, resulting in hallucinated or incorrect function calls and suboptimal parameter choices. Moreover, the use of RAG significantly increased token usage and inference time, raising concerns about scalability and cost efficiency.
Evaluation Metric Limitations.
Our automated metrics for Accuracy & Depth (Spearman $\rho=0.14$), Logical Flow ($\rho=0.16$), and Visual Consistency ($\rho=0.17$) show weak correlation with human ratings, while Visual Relevance ($\rho=0.72$) and Element Layout ($\rho=0.42$) achieve moderate to high agreement with good p-values (0.001 and 0.03 respectively) to show statistical significance. We acknowledge that our automated metrics have room for improvement, particularly for aspects like narrative coherence and the nuances of visual consistency. Current video understanding capabilities in Vision Language Models (VLMs) are still in early stages of development. Future work could explore fine-tuning specialized video understanding models (e.g., VideoScore He et al. (2024b)) for assessing these types of tasks more accurately. We agree that visual consistency is part of the visual assessment and will refine the evaluation design in future iterations.
Code Generation Fragility.
The system remains vulnerable to common coding errors, including LaTeX rendering failures, incorrect function calls, and general Python errors such as undefined variables and missing imports. These issues were detailed in Section 4.3 and highlight the fragile nature of current code-generation approaches. Despite the implementation of a retry mechanism with up to five attempts, code reliability remain critical challenges.
@@ Appendix A Gallery
Appendix B Runtime Statistics
Appendix C Preference Study
To complement our automated metrics, we conducted a human evaluation study. 18 annotators were asked to rank videos on 3 topics generated by three different models (o3-mini, Gemini 2.0 Flash, and GPT-4o) for the same set of theorems, without revealing the model names. Annotators evaluated the videos based on clarity and visual appeal. The results, summarized in Table 6, indicate that videos generated by Gemini 2.0 Flash were most frequently ranked highest for both clarity and visual appeal. o3-mini generated videos were preferred over those from GPT-4o. This human preference study provides complementary insights into the perceived quality of the generated videos.
Appendix D Implementation Details
D.1 Human Annotation Process
We recruited 12 student volunteers in our annotation process for the human metric. We explained to the annotators that their annotations were to be used in our study only and would not be released publicly.
D.2 Technical Implementation
To aid reproducibility, we provide the following key implementation details. Manim Versions:
  • Manim Community Edition version 0.18.1,
  • ManimPango 0.6.0,
  • manim-physics 0.4.0,
  • manim-ml Helbling et al. (2023) 0.0.24,
  • manim-chemistry 0.4.4,
  • manim-dsa Missagia (2025) 0.2.0,
  • manim-circuit 0.0.3
RAG System: We utilize ChromaDB as the vector store. Documentation (Markdown and Python files from Manim core and specified plugins) is chunked using Langchain’s RecursiveCharacterTextSplitter with default settings for the respective languages. The embedding model used is text-embedding-005 from Google Vertex AI.
RAG Retrieval Threshold: A relevance score threshold of 0.5 is used during similarity search in both core and plugin vector stores. We retrieve k=2 documents per query by default.
D.3 Prompt Templates
@@
D.4 Computational Resources and Costs
The primary LLM computations were performed using closed-source model APIs. The average number of tokens, cost, and inference time per video generation are detailed in Table 5. Local Manim rendering does not require a GPU for the types of animations generated.
Appendix E Potentials for Future Research
Appendix F User Feedback on Educational Use
We conducted a small-scale user study with 15 students and 8 teachers at universities to gather feedback on the educational clarity and engagement of AI-generated videos on easy topics. Participants rated videos on two topics (Bubble Sort and Hamming Distance) on a scale of 1 to 5 for clarity and engagement. As shown in Table 7, the mean scores for clarity and engagement were high. Over 90% of participants expressed interest in using such videos in a classroom setting, although the majority also believed there is room for content improvement.
Appendix G Definition of Agentic
Our use of the term “agentic" follows the definition Russell and Norvig (2020): “An agent is anything that can be viewed as perceiving its environment through sensors and acting upon that environment through actuators." In the context of TheoremExplainAgent, the language model acts as an agent that perceives inputs (such as the theorem description and feedback from code execution errors) and acts upon its environment by generating Manim code to create video scenes.
Agent Input Tokens Output Tokens Cost(USD) Time(s)
GPT-4o 350000 84000 1.71 1120
GPT-4o + RAG 840000 84000 2.94 1260
Claude 3.5-Sonnet v1 350000 91000 2.42 2240
Claude 3.5-Sonnet v1 + RAG 1050000 101500 4.67 2380
Gemini 2.0-Flash 595000 119000 0.1 1120
Gemini 2.0-Flash + RAG 1120000 119000 0.16 1260
o3-mini (medium) 434000 154000 1.16 1680
o3-mini (medium) + RAG 945000 154000 1.72 1820
Model Agent Clarity Top-Rank % Visual Top-Rank %
Gemini-2.0-Flash Gemini-Team et al. (2024) 70.6% 61.8%
o3-mini OpenAI (2025) 20.6% 23.5%
GPT-4o OpenAI (2023) 8.8% 14.7%
Table 6: Preference Study: Percentage of times each model was ranked highest by human annotators for video clarity and visual appeal.
Metric Topic 1: Bubble Sort Topic 2: Hamming Distance
Clarity (Mean $\pm$ SD) 4.13 $\pm$ 0.87 4.35 $\pm$ 0.83
Engagement (Mean $\pm$ SD) 3.57 $\pm$ 1.16 3.78 $\pm$ 1.09
Would recommend for classroom use 13% 35%
Would recommend for classroom use if video content improved 87% 61%
Would not recommend for classroom use 0% 4%
Table 7: User Feedback on Educational Usefulness.